OpenStudy (anonymous):

solve for x 3x+5.4=9.6

OpenStudy (anonymous):

(9.6-5.4)/3 will give you x

OpenStudy (anonymous):

subtract 5.4 from both sides 3x + 5.4 = 9.6 - 5.4 -5.4 3x = 4.3 Now divide /3 /3 x = 1.4

OpenStudy (geometry_hater):

Subtract 9.6 by 5.4 giving you something Then divide by 3 and that will be your x

OpenStudy (anonymous):

you need to put all your numbers on one side of the equation and your x's on the other side of the equation subtract your numbers divide each side by the constant next to the varible to get x by itsself then you have your answer!